Tohban Report 2017-06-21
Tohban Reports | |
---|---|
Start Date: | 14 Jun 2017 |
End Date: | 21 Jun 2017 |
Tohban: | Pascal Saint-Hilaire |
Tohban email: | pascal@ssl.berkeley.edu |
Next Tohban: | TBD |
List all reports |
Solar Activity
Solar activity during last week was low and currently is at low levels. GOES baseline in the high As. A new AR on the eastern limb has low probabilities for C-flares.
How many GOES flares occurred?
Flares above B, C, M, X class were 15 0 0 0
And how many of these are listed in the RHESSI flare list?
Flares above B, C, M, X class were 8 0 0 0
And how many had EXCELLENT coverage?
Flares above B, C, M, X class were 8 0 0 0
There were RHESSI flares/GOES flares 16 / 15 over the time range 13-Jun-17 20-Jun-17
Memory Management
The SSR is below 30% at the end of the daily pass set. TBC
Spacecraft Status
Detectors 1, 3, 6 and 8 are opertaing and collecting science data. Cold plate temperatures are and K. Cold tip temperatures are and K.
The Cryocooler is operating with a power of ~74.2 W and the acceleration is at ~10.8mg (? it was 14.8 last week???)
Data Gaps
No data gaps of note from the 12th to the 19th.
Detector issues
Detector front livetimes seem to vary by 4-5% over the course of a day or so. The weaker, G1, has ~70/75% min/max.
Other notes
Spacecraft is in good conditions and operating.
Spacecraft Management
Decimation | Active/Vigorous |
HLAT Decimation | Rear decimation |
Night time data (fronts) | plus/minus 4 minutes from daylight |
Night time data (rears) | plus/minus 4 minutes from daylight |
Require extra passes? | No |
Requirement for moving pointer? | No |
Attenuator operation | logic enabled |
Detector problems? | No problems. |